OpenStudy (anonymous):

y varies inversely as the square of x and the cube root of z, but inversely as the square root of w. variation equation if y=7 when x=3, z=27, and w=33

hartnn (hartnn):

y varies directly as the square of x ----> \(y =k_1 x^2\) y varies directly as the cube root of z---->\(y=k_2z^3\) inversely as the square root of w -----> \(y=k_3/\sqrt w\) combining all these 3 . \(y=kx^2z^3/\sqrt w\) gott his ?

OpenStudy (johnweldon1993):

You mean \[y = k_2\sqrt[3]{z}\] right?

hartnn (hartnn):

oops, i missed the "root" part ... y varies directly as the cube root of z---->\(y=k_2\sqrt[3]z\) so, \(y=kx^2\sqrt[3]z/\sqrt w\)

OpenStudy (anonymous):

I'm with you so far. But don't I know have to solve for K, and plug that into my original equation?

hartnn (hartnn):

yes, plug in y=7 when x=2, z=3, and w=25 in that equation!

OpenStudy (anonymous):

And that's where it gets really funky for me.

hartnn (hartnn):

ok, then take it one by one x= 2, x^2 =.. ?

OpenStudy (anonymous):

\[35 = 4K \sqrt[3]{3}\]

OpenStudy (anonymous):

Is as far as I can seem to get.

hartnn (hartnn):

ok, correct, now just isolate k divide \(4\sqrt[3]3\) on both sides

hartnn (hartnn):

if you need to rationalize the denominator, then use, \(\dfrac{1}{\sqrt[3]3}=\dfrac{1}{3^{1/3}}=\dfrac{3^{-1/3}}{1}\)

OpenStudy (anonymous):

Reviewing my notes for rationalizing denominators. BRB

hartnn (hartnn):

i've used this rule \(\large \dfrac{1}{x^n}=x^{-n}\)

OpenStudy (anonymous):

um. I got \[\frac{ 140\sqrt[3]{3} }{ 48 }\]

OpenStudy (anonymous):

Multiplied conjugates?

hartnn (hartnn):

do you have choices ? then post them..

OpenStudy (anonymous):

woops, 3 was supposed to be Z

OpenStudy (anonymous):

Is that correct? No choices.

OpenStudy (anonymous):

Ok thanks!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

i missed the opportunity to welcome you here, sin you are new,so, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

go back to your recent question...